Let V and W be two vector spaces over the field F and T, T1, T2 be linear transformations from V to W. Prove the following:

(a) rank(αT) = rank(T), for all α ∈ F and α 6= 0.

(b) |rank(T1) − rank(T2)| ≤ rank(T1 + T2) ≤ rank(T1) + rank(T2


1
Expert's answer
2021-07-07T07:27:39-0400
  • For "\alpha\neq 0" we have "\text{Range} (T)=\text{Range}(\alpha T)", as for any "y\in \text{Range}(T)" there is "x\in V" such that "Tx=y". Now as "F" is a field, "\alpha^{-1}" exists and thus "\alpha T(\alpha^{-1}x)=\alpha \alpha^{-1}Tx=y" by linearity of "T" and thus "y\in \text{Range}(\alpha T)" and "\text{Range}(T)\subseteq \text{Range}(\alpha T)" . We can apply the same argument to the other inclusion to conclude that "\text{Range}(T)=\text{Range}(\alpha T)" and thus "rg(T)=rg (\alpha T)".
  • Let us first prove the right inequality. We have "(T_1+T_2)(x)=T_1(x)+T_2(x)" and thus "\text{Range}(T_1+T_2)\subseteq \text{Range}(T_1)+\text{Range}(T_2)", where plus on the right is the plus in the sense of vector subspaces of "W". As "\dim(A+B)\leq \dim(A)+\dim(B)" for "A,B" vector spaces, we conclude that "rg(T_1+T_2)\leq \dim(\text{Range}(T_1)+\text{Range}(T_2))\leq rg(T_1)+rg(T_2)".

Now to find the left inequality, we just need to use the inequality we just found by replacing "T_2" by "-T_2" and "T_1" by "T_1+T_2" : "rg(T_1+T_2-T_2)\leq rg(T_1+T_2)+rg(-T_2)". Now as "-1\neq 0" in "F", we have "rg(-T_2)=rg(T_2)" and therefore "rg(T_1)\leq rg(T_1+T_2)+rg(T_2)" or we can rewrite it as "rg(T_1)-rg(T_2)\leq rg(T_1+T_2)". Applying the same symmetric argument to "-T_1" and "T_1+T_2" we find that "rg(T_2)-rg(T_1)\leq rg(T_1+T_2)". Combining the two inequalities we conclude that "|rg(T_1)-rg(T_2)|\leq rg(T_1+T_2)".